Four items are selected at a random from a box of 12 items and inspected the box is rejected if more than 1 item is found to be faulty.if there are 3 faulty items in a box find the probability that the box is accepted
1
Expert's answer
2020-08-10T18:40:18-0400

We can use combination ("C^n_k" ) which shows how may ways to pick n items from a collection of k items, when the order of selection does not matter. "C^k_n=n!\/(k!(n-k)!)".

We can choose 4 items from 12 in "C^4_{12}=12!\/(4!*8!)=495" ways. This is a total number of different variants.

Now we should find how many ways to pick 4 good items or 3 good items and 1 faulty item and sum them:

We can choose 4 good items in "C^4_{9}=9!\/(4!*5!)=126" ways (because we have only 9 good items in the box and we pick 4 of them).

We can choose 3 good items and 1 faulty item in in "C^3_9*C^1_3=9!\/(3!*6!)*3!\/(1!*2!)=252" ways (because we picks 3 good items out of 9 and 1 faulty item out of 3).

Thus, the probability that the box is accepted:

"(C^4_{9}+C^3_9*C^1_3)\/C^4_{12}=(126+252)\/495=0.7636"